OpenStudy (anonymous):

PLEASE HELP a dice is rolled 8 times. find the probability. p(getting at least 7 even numbers) a. 1/128 b. 7/32 c. 1/32 d. 1/16

OpenStudy (anonymous):

my explanation was that there are 48 outcomes and there are 3 even nos. that we want 7 times so 7 x 3 = 21 Final ans = 21/48

OpenStudy (anonymous):

obviously it is wrong pls explain where i am going wrong

OpenStudy (amistre64):

at least 7 means, P(7) or P(8) or P(9) or .... so, what is the probability of getting 7 even numbers ... what is the probability of getting 8 even numbers? then add the results

OpenStudy (amistre64):

do you know how to do i binomial probability?\[\binom{n}{k}p^kq^{n-k }\]

OpenStudy (anonymous):

no

OpenStudy (amistre64):

hmmm, well do you know the probability of getting an even or odd number on 1 roll of the dice?

OpenStudy (anonymous):

.5 ?

OpenStudy (amistre64):

yes

OpenStudy (amistre64):

now, if we setup an 8 roll situation we get: eeeeeeeo, 7 even and 1 odd how many ways are there to permute this arrangement? in other words, how many ways can we resort this?

OpenStudy (amistre64):

1 eeeeeeeo 2 eeeeeeoe 3 eeeeeoee 4 eeeeoeee 5 eeeoeeee 6 eeoeeeee 7 eoeeeeee 8 oeeeeeee i count 8 ways that we can arrange the setup. so 8 times, .5^7 * .5^1 this gives us the probability of rolling 7 even numbers out of 8 rolls

OpenStudy (anonymous):

so 8 times, .5^7 * .5^1 ? this part makes no sense

OpenStudy (amistre64):

do you agree that there are 8 ways to arrange how the outcome of the dice happen in order to get us 7 even numbers and 1 odd number?

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

all thats left to do is determine the probability of getting: an even, and an even, and an even, and an even, and an even, and an even, and an even, and an odd. P(even) and P(even) and P(even) and P(even) and P(even) and P(even) and P(even) and P(odd) or written another way P(even) * P(even) *P(even) *P(even) *P(even) *P(even) *P(even) *P(odd) or better yet, since P(even) = P(odd) = .5 .5 * .5 * .5 * .5 * .5 * .5 * .5 * .5

OpenStudy (anonymous):

ok

OpenStudy (amistre64):

so, 8 ways to get: .5^7 * .5^1 7 even and 1 odd

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

P(7even and 1 odd): 8*.5^8 and there is only 1 way to get 8 even so P(8even and 0odd) = 1*.5^8 the probability of getting one or the other is just the sum of them together 8*.5^8 + .5^8. but since your options are in fractions it might be best to use 1/2 instead of .5

OpenStudy (anonymous):

so what is the final ans?

OpenStudy (anonymous):

(1/2)^8

OpenStudy (anonymous):

@amistre64

OpenStudy (amistre64):

the final answer is: 8*.5^8 + .5^8, simplify to your hearts content

OpenStudy (amistre64):

8/2^8 + 1/2^8 = 9/2^8

OpenStudy (anonymous):

C ? or B ?

OpenStudy (amistre64):

which one is equal to 9/2^8?

OpenStudy (amistre64):

since 9 and 2^8 have no common factors .... id pick the one with a 9 up top :/

OpenStudy (anonymous):

which choice is it now??

OpenStudy (anonymous):

i wrote my choices incorrectly last time

OpenStudy (amistre64):

... sigh. why did you change them?